# Uncertainty-Aware Explanations Through Probabilistic Self-Explainable Neural Networks

### v1

@@ Abstract
The lack of transparency of Deep Neural Networks continues to be a limitation that severely undermines their reliability and usage in high-stakes applications. Promising approaches to overcome such limitations are Prototype-Based Self-Explainable Neural Networks (PSENNs), whose predictions rely on the similarity between the input at hand and a set of prototypical representations of the output classes, offering therefore a deep, yet transparent-by-design, architecture. So far, such models have been designed by considering pointwise estimates for the prototypes, which remain fixed after the learning phase of the model. In this paper, we introduce a probabilistic reformulation of PSENNs, called Prob-PSENN, which replaces point estimates for the prototypes with probability distributions over their values. This provides not only a more flexible framework for an end-to-end learning of prototypes, but can also capture the explanatory uncertainty of the model, which is a missing feature in previous approaches. In addition, since the prototypes determine both the explanation and the prediction, Prob-PSENNs allow us to detect when the model is making uninformed or uncertain predictions, and to obtain valid explanations for them. Our experiments demonstrate that Prob-PSENNs provide more meaningful and robust explanations than their non-probabilistic counterparts, thus enhancing the explainability and reliability of the models.
@@ 1 Introduction
The complexity in the architectures of state-of-the-art Deep Neural Networks (DNNs) largely accounts for their “black box” nature Rudin (2019), which is in conflict with one of the basic requirements for trustworthiness Ashoori and Weisz (2019): to be able to explain and understand the decisions of the model. Although several strategies have been proposed in order to explain black box models Zhang et al. (2021), they often provide only partial information about their inner workings, such as which parts of the input at hand most condition the output Simonyan et al. (2014); Selvaraju et al. (2017); Sundararajan et al. (2017); Bach et al. (2015); Shrikumar et al. (2017); Springenberg et al. (2015). However, these approaches do not provide information about how the model processes that information or why they imply the output Rudin (2019); Hase et al. (2019); Adebayo et al. (2018); Lipton (2018). In order to achieve a more transparent classification process, recent works advocate the use of Prototype-Based Self-Explainable Neural Networks (PSENNs, for simplicity) (Alvarez-Melis and Jaakkola, 2018; Chen et al., 2019; Hase et al., 2019; Li et al., 2018; Gautam et al., 2022), which are trained to jointly maximize their prediction performance and their explainability. These approaches rely primarily on prototype-based models, in which the prototypes are learned during the training phase, aiming to capture discriminative and also semantically-meaningful features. In this way, the output classification is based on the similarity between the input and the learned prototypes, providing an intuitive and human-understandable classification process.
@@
On the other hand, current DNNs models, including the PSENNs discussed above, are optimized with the aim of finding the point estimate values of the parameters that minimize a given loss term, leading to models with a deterministic inference process. An important drawback of these approaches is that, generally, it is not possible to measure to what extent the model is confident in its own prediction, known in the literature as the epistemic uncertainty of the model Hüllermeier and Waegeman (2021); Kwon et al. (2022); Gal (2016). Indeed, it has been shown that DNNs tend to be overconfident in their predictions Guo et al. (2017), classifying inputs with high probability even when random or out-of-distribution inputs are provided. Being unable to detect when the model is making such unreliable or uninformed decisions dramatically reduces the trustworthiness and reliability of the model, and is, therefore, of great concern in high-stakes applications. To address this issue, recent works proposed replacing point estimates for the parameters of the model with a probability distribution over their values Goan and Fookes (2020). This leads to a stochastic model, for which it is possible to measure the aforementioned types of uncertainty, and, thus, identify when the model is making uncertain predictions that should not be trusted.
@@
In this paper, we aim to address the aforementioned drawbacks and challenges by leveraging recent tools from uncertainty estimation in DNNs, and combining them with PSENN architectures in order to develop Probabilistic Self-Explainable Neural Networks (Prob-PSENNs). The key feature of this model is that the prototypes are defined as random variables, for which suitable probability distributions over their values are learned during the training phase of the model – jointly with the rest of the parameters of the network. This probabilistic paradigm allows us not only to improve the explainability of the classification process, but also to capture different sources of uncertainty regarding the explanation. Furthermore, since the output of the PSENNs directly depends on the explanatory component Li et al. (2018), we can establish connections between their corresponding uncertainties, enabling a more thorough analysis of the prediction process, and allowing the models to self-explain their own uncertainties.
@@
    First, we introduce Prob-PSENN, a paradigm shift for PSENNs, which relies on probabilistic modeling of the explanatory component, thus replacing point estimates for the prototypes with probability distributions over their values.
@@
    Furthermore, Prob-PSENNs allow us, for the first time, to model the uncertainty in the explanations and in the predictions, which is a missing feature in conventional PSENNs. More specifically, we formalize different explanatory uncertainty notions, and show how Prob-PSENNs enable not only the quantification of those uncertainties, but also their explanations, thus increasing reliability and trustworthiness.
@@
    The effectiveness of Prob-PSENN is evaluated in several classification tasks, demonstrating that it can effectively learn highly diverse and prototypical representations of the output classes, capture the model uncertainties and maintain competitive predictive performance.
@@ 2 Preliminaries: Prototype-Based Self-Explainable Neural Networks (PSENNs)
Let us consider a classification problem in which the goal is to classify $d$-dimensional inputs $x\in\mathbb{R}^{d}$ into one of the $c$ possible classes in $Y=\{y_{1},y_{2},\dots,y_{c}\}$. Using as baseline the architecture introduced in Li et al. (2018), illustrated in Figure 1, the main components of the PSENNs are described as follows. We refer the reader to Appendix A for a justification for the choice of this baseline, as well as for a discussion on related works.
Figure 1: Architecture of the Prototype-Based Self-Explainable Neural Network Li et al. (2018) described in Section 2, illustrated for a handwritten digit classification task Lecun et al. (1998).
@@
Taking $\delta_{x}$ as the input, the output provided by the model will be determined by a prototype classifier $h:\mathbb{R}^{m}\rightarrow[0,1]^{c}$. As in Li et al. (2018), a classification network $h(\delta_{x})=s(W\delta_{x})$ will be assumed, being $W$ a $c\times m$ weight matrix and $s(\cdot)$ the softmax function. For simplicity, the end-to-end classification process will be denoted as $f(x)=h(\delta_{x})$. Notice that $f(x)$ outputs a vector representing the probability with which the input belongs to each of the possible classes of the problem, according to the model. The probability assigned to each class $y_{i}$ will be denoted as $f_{i}(x)$, $1\leq i\leq c$, and thus we can define the output as: $f(x)=\bigl{(}f_{1}(x),f_{2}(x),\dots,f_{c}(x)\bigr{)}$. Generally, the class with the highest probability is considered the output class of the prediction, what will be denoted as $f_{*}(x)=\operatorname*{arg\,max}_{y_{i}\in Y}f_{i}(x)$.
@@
Since the output class is based on the distance-vector $\delta_{x}$, the explanation $\xi_{x}$ for the prediction $f(x)$ can be determined also by the similarity between the prototypes and the input at hand: $\xi_{x}=\bigl{\{}\{g(r_{1}),\delta_{x,1}\},\{g(r_{2}),\delta_{x,2}\},\dots,\{g(r_{m}),\delta_{x,m}\}\bigr{\}},$ being $g(\cdot)$ the decoder defined above. Notice that the interpretability of the model relies, therefore, on to what extent the decoded prototypes are capable of capturing relevant and semantically-meaningful features representing each class of the problem, which is enforced during the training phase by means of including interpretability regularizers in the loss function Li et al. (2018).
3 Probabilistic Self-Explainable Neural Networks (Prob-PSENNs)
As mentioned before, PSENNs treat the prototypes as network parameters, for which point estimate values are computed and hence are fixed once the training is completed. Furthermore, the number of prototypes $m$ needs to be finite and defined beforehand, even if, in practice, it might result cumbersome or infeasible to determine how many prototypes will be required to capture sufficiently diverse characterizations of the classes.
To address these issues, we introduce Prob-PSENNs, generalizing the prototypes by means of defining a probability distribution over their values. The main motivation for this choice is twofold. On the one hand, for each class of the problem, there might exist several sets of prototypes $R$ capable of representing the data with a comparable effectiveness. For illustration, let us consider a handwritten digit classification scenario, and the fact that there are several ways of representing each digit. This implies that there are several representative prototypes for each class. Consequently, instead of relying on a single set of prototypes, it might be more reasonable to consider all those possible variations in order to classify an input. This probabilistic reformulation also enables a flexible way to decide the variety of the prototypes required to represent the classes, which will be determined, assuming one prototype per class, by the variance of the distribution over the prototypes. On the other hand, the choice of the prototypes in $R$ conditions not only the output but also the explanation. Therefore, placing a probability distribution over the prototypes allows us to enable a probabilistic framework for the generation of the explanation as well, which can be employed to capture the uncertainty in the explanations of the model, and, thereby, produce more informative and useful explanations.
@@ 3.1 Predictive distribution
Instead of considering a point estimate for $R$, we will consider it a random variable, for which we will use the notation $\bm{R}$. Furthermore, we will assume $m=c$ prototypes, such that $\bm{r}_{i}$ represents the class $y_{i}$, $1\leq i\leq c$. Due to the randomness in the network’s prototypes, the inference of the model will be stochastic, and will be denoted as $f^{\bm{R}}(x)$. In this way, a predictive distribution for the output can be defined:
@@ 3.2 Capturing the predictive uncertainty
The predictive distribution defined in Equation (1) allows us to better model the predictive uncertainty in comparison to its deterministic counterpart. First, the total predictive uncertainty can be measured by means of the entropy Shannon (1948) of the prediction: ${H[p(y|x)]=-\sum_{i=1}^{c}p(y=y_{i}|x)\cdot\log p(y=y_{i}|x)}$· However, for a model with fixed (pointwise) values for the parameters, the source of this uncertainty will be only aleatoric uncertainty, which refers to the possibly unavoidable noise, randomness or ambiguity inherent in the data. Considering a probability distribution for the parameter values instead of point estimates allows us to capture also the epistemic uncertainty Hüllermeier and Waegeman (2021); Kendall and Gal (2017); Kwon et al. (2022), which accounts for the model uncertainty about the true parameters modeling the data (e.g., as a consequence of insufficient training data). This enables a safer use of the model in critical tasks, since it allows us to identify when the model is making unreliable predictions. An illustrative comparison between aleatoric and epistemic uncertainty is provided in Figure 2 (left).
Figure 2: Comparison between the predictive and explanatory uncertainty, illustrated for a binary classification problem with a two-dimensional input space.
Thus, for the predictive distribution defined in Equation (1), $H[p(y|x)]$ encapsulates both aleatoric and epistemic uncertainty. To measure epistemic uncertainty only, the mutual information between the output $\bm{y}$ and the parameters $\bm{R}$ can be employed Houlsby et al. (2011); Gal (2016); Hüllermeier and Waegeman (2021); Depeweg et al. (2018), which, in our case, can be formalized as: $I(\bm{y},\bm{R}|x)=H(\bm{y}|x)-\mathbb{E}_{p(\bm{R})}[H(\bm{y}|x,\bm{R})],$ while $\mathbb{E}_{p(\bm{R})}[H(\bm{y}|x,\bm{R})]$ represents a measure of aleatoric uncertainty. Notice that the values of all these metrics can be normalized to the range $[0,1]$, where $1$ would represents the maximum uncertainty, by dividing the resulting value by the entropy of the uniform categorical distribution of $c$ categories. In practice, $H[p(y|x)]$ can be approximated by $H[\tilde{p}(y|x)]$ (see Equation 2).
@@ 3.3 Capturing the explanation uncertainty
With a PSENN with fixed prototypes, only a point estimate $\delta_{x}$ can be provided for the input at hand, that is, a point estimate on how close the input is to the particular prototypes selected to represent each class, as illustrated in Figure 3-(a). In contrast, the probability distribution over the prototypes considered in Prob-PSENNs induces a probability distribution over each value of $\delta_{x}$, $\,p(\bm{\delta}_{x,i})\!=\!p\bigl{(}\phi(e_{x},\bm{r}_{i})\bigr{)}$, as exemplified in Figure 3-(b) and Figure 3-(c). This allows us to compute not only more robust and representative information for each value of $\delta_{x}$, but also to compute the uncertainty in the explanation. Furthermore, this enables a direct and explicit connection between the explanation uncertainty and the output uncertainty, being therefore a key advantage of Prob-PSENNs.
Types of explanation uncertainty
Similarly to the predictive uncertainty, different types of uncertainty can be attributed to the explanations of Prob-PSENNs. In the context of explanations, we will denominate aleatoric uncertainty to the uncertainty in determining which class-prototypes best represent the input at hand, that is, in determining which prototypes are closest to the encoded input, possibly due to ambiguity in the representation of the input. Therefore, high aleatoric uncertainty will imply highly overlapped distance-distributions $p(\bm{\delta}_{x,i})$. In contrast, epistemic explanatory uncertainty will be high for those inputs that lie far from the training distribution, and, consequently, also far from the distribution over the prototypes, resulting in uninformative and non-representative explanations for that input. Therefore, high epistemic uncertainty will imply large distances with respect to all the prototypes. An illustrative representation of these types of explanatory uncertainty is provided in Figure 2 (right).
Quantifying explanation uncertainty
Given that, in the general case, computing $p(\bm{\delta}_{x,i})$ will be intractable both analytically and numerically, we will rely on sample-based estimators in order to compute meaningful information about those distributions. To begin with, being $\delta_{x}^{(1)},\dots,\delta_{x}^{(N)}$ the distance-vectors obtained for $N$ inferences of the model (i.e., for $N$ random sets of prototypes), the sample mean ${\bar{\delta}_{x,i}=\frac{1}{N}\sum_{n=1}^{N}\delta_{x,i}^{(n)}}$ can inform us about the average similarity between the input and the prototypes of the class $y_{i}$, while the sample variance Var ${\bigl{(}p(\bm{\delta}_{x,i})\bigr{)}\approx\frac{1}{N-1}\sum_{n=1}^{N}\bigl{(}\delta_{x,i}^{(n)}-\bar{\delta}_{x,i}\bigr{)}^{2}}$ can be considered a measure of uncertainty about those similarities.

### v2

@@ Abstract
The lack of transparency of Deep Neural Networks continues to be a limitation that severely undermines their reliability and usage in high-stakes applications. Promising approaches to overcome such limitations are Prototype-Based Self-Explainable Neural Networks (PSENNs), whose predictions rely on the similarity between the input at hand and a set of prototypical representations of the output classes, offering therefore a deep, yet transparent-by-design, architecture. In this paper, we introduce a probabilistic reformulation of PSENNs, called Prob-PSENN, which replaces point estimates for the prototypes with probability distributions over their values. This provides not only a more flexible framework for an end-to-end learning of prototypes, but can also capture the explanatory uncertainty of the model, which is a missing feature in previous approaches. In addition, since the prototypes determine both the explanation and the prediction, Prob-PSENNs allow us to detect when the model is making uninformed or uncertain predictions, and to obtain valid explanations for them. Our experiments demonstrate that Prob-PSENNs provide more meaningful and robust explanations than their non-probabilistic counterparts, while remaining competitive in terms of predictive performance, thus enhancing the explainability and reliability of the models.
@@ 1 Introduction
The complexity in the architectures of state-of-the-art Deep Neural Networks (DNNs) largely accounts for their “black box” nature [34], which is in conflict with one of the basic requirements for trustworthiness [4]: to be able to explain and understand the decisions of the model. Although several strategies have been proposed in order to explain black box models [46], they often provide only partial information about their inner workings, such as which parts of the input at hand most condition the output [38, 35, 40, 5, 37, 39]. However, these approaches do not provide information about how the model processes that information or why they imply the output [34, 17, 1, 28]. In order to achieve a more transparent classification process, recent works advocate the use of Prototype-Based Self-Explainable Neural Networks (PSENNs, for simplicity) [3, 7, 17, 27, 13], which are trained to jointly maximize their prediction performance and their explainability. These approaches rely primarily on prototype-based models, in which the prototypes are learned during the training phase, aiming to capture discriminative and also semantically-meaningful features. In this way, the output classification is based on the similarity between the input and the learned prototypes, providing an intuitive and human-understandable classification process.
@@
On the other hand, current DNNs, including the PSENNs discussed above, are generally optimized with the aim of finding the point estimate values of the parameters that minimize a given loss term, leading to models with a deterministic inference process. An important drawback of these approaches is that, generally, it is not possible to measure to what extent the model is confident in its own prediction, known in the literature as the epistemic uncertainty of the model [19, 24, 12]. Indeed, it has been shown that DNNs tend to be overconfident in their predictions [16], classifying inputs with high probability even when random or out-of-distribution inputs are provided. Being unable to detect when the model is making such unreliable or uninformed decisions dramatically reduces the trustworthiness and reliability of the model, and is, therefore, of great concern in high-stakes applications, as recognized in the EU AI Act [11]. To address this issue, recent works proposed replacing point estimates for the parameters of the model with a probability distribution over their values [15]. This leads to a stochastic model, for which it is possible to measure the aforementioned types of uncertainty, and, thus, identify when the model is making uncertain predictions that should not be trusted.
@@
In this paper, we aim to address the aforementioned drawbacks and challenges by leveraging recent tools from uncertainty estimation in DNNs, and combining them with PSENN architectures in order to develop Probabilistic PSENNs (Prob-PSENNs). The key feature of this model is that the prototypes are defined as random variables, for which suitable probability distributions are learned during the training phase of the model – jointly with the rest of the parameters of the network. This probabilistic paradigm allows us not only to improve the explainability of the classification process, but also to capture different sources of uncertainty regarding both the output and the explanation. Furthermore, since the output of the PSENNs directly depends on the explanatory component [27, 7, 13], we can establish connections between their corresponding uncertainties, enabling a more thorough analysis of the prediction process, and allowing the models to self-explain their own uncertainties.
@@
    We introduce Prob-PSENN, a probabilistic reformulation of PSENNs that enhances trustworthiness by addressing two critical aspects: (i) interpretability and transparency, and (ii) uncertainty quantification.
@@
    Prob-PSENNs allow us, for the first time, to model the uncertainty in both the explanations and predictions, which is a missing feature in conventional PSENNs. We formalize different explanatory uncertainty notions, and show how Prob-PSENNs can (i) quantify those uncertainties, (ii) explain them, and (iii) leverage them to improve their predictive performance.
@@
    The effectiveness of Prob-PSENNs is evaluated in several classification tasks, demonstrating their reliability and trustworthiness.
@@ 2 Preliminaries: Prototype-Based Self-Explainable Neural Networks (PSENNs)
Let us consider a classification problem in which the goal is to classify $d$-dimensional inputs $x\in\mathbb{R}^{d}$ into one of the $c$ possible classes in $Y=\{y_{1},y_{2},\dots,y_{c}\}$. Building on the frameworks introduced in [27, 13], the core components of the PSENNs are described below. We refer the reader to Appendix A for a justification of the choice of these frameworks, as well as a discussion of related works.
Figure 1: Architecture of the proposed Probabilistic PSENN (Prob-PSENN). The model replaces deterministic prototypes with probabilistic prototype distributions, enabling the sampling of diverse yet representative prototypes from the latent space. To account for prototype variability, $N$ samples of the prototypes can be considered during the inference process, producing diverse predictions that enable capturing both explanatory and predictive uncertainties.
@@
Taking $\delta_{x}$ as the input, the output provided by the model will be determined by a prototype classifier $h:\mathbb{R}^{m}\rightarrow[0,1]^{c}$. As in [27], a classification network $h(\delta_{x})=s(W\delta_{x})$ will be assumed, where $W$ is a $c\times m$ weight matrix and $s(\cdot)$ the softmax function. For simplicity, the end-to-end classification process will be denoted as $f(x)=h(\delta_{x})$. Notice that $f(x)$ outputs a vector representing the probability with which the input belongs to each of the possible classes of the problem, according to the model. The probability assigned to each class $y_{i}$ will be denoted as $f_{i}(x)$, $1\leq i\leq c$, and thus we can define the output as: $f(x)=\bigl{(}f_{1}(x),f_{2}(x),\dots,f_{c}(x)\bigr{)}$.
@@
Since the output class is based on the distance-vector $\delta_{x}$, the explanation $\xi_{x}$ for the prediction $f(x)$ can be determined also by the similarity between the prototypes and the input at hand: $\xi_{x}=\bigl{\{}\{g(r_{1}),\delta_{x,1}\},\{g(r_{2}),\delta_{x,2}\},\dots,\{g(r_{m}),\delta_{x,m}\}\bigr{\}},$ where $g(\cdot)$ is the decoder defined above. The interpretability of the model relies, therefore, on to what extent the decoded prototypes are capable of capturing relevant and semantically-meaningful features representing each class of the problem, which is enforced during the training phase by means of including interpretability regularizers in the loss function [27, 13].
3 Probabilistic Prototype-Based Self-Explainable Neural Networks (Prob-PSENNs)
As mentioned above, PSENNs treat the prototypes as network parameters, for which point estimate values are computed and hence are fixed once the training is completed. In this section, we introduce Prob-PSENNs, which generalize the prototypes by means of defining and learning a probability distribution over their values, as we illustrate in Figure 1. The main motivations for this probabilistic formulation are discussed below.
First, for each class of the problem, there might exist several sets of prototypes $R$ capable of representing the data with comparable effectiveness. For illustration, let us consider a handwritten digit classification scenario, and the fact that there are several ways of representing each digit. This implies that there are several representative prototypes for each class. Consequently, instead of relying on a single set of prototypes, it might be more reasonable to consider all those possible variations in order to classify an input. Furthermore, in previous approaches, the number of prototypes $m$ needs to be finite and defined beforehand, even if, in practice, it might be cumbersome or infeasible to determine how many prototypes will be required to capture sufficiently diverse characterizations of the classes. In contrast, our probabilistic reformulation enables a more flexible way to decide the variety of the prototypes required to represent the classes, which will be determined, assuming one prototype distribution per class, by the variance of the distributions.
Finally, the choice of the prototypes in $R$ conditions not only the output but also the explanation. Therefore, placing a probability distribution over the prototypes allows us to enable a novel probabilistic framework for the generation of the explanation as well, which can be employed to capture different sources of uncertainty in the explanations of the model, and thereby, produce more informative and useful explanations.
@@ 3.1 Predictive distribution
Instead of considering a point estimate for $R$, we will use a random variable, denoted $\bm{R}$. In this way, we can consider $m=c$ random class-prototypes, such that $\bm{r}_{i}$ represents the class $y_{i}$, $1\leq i\leq c$. This eliminates the need to estimate $m$ and avoids having to restrict the model to a fixed number of prototypes per class, as each $\bm{r}_{i}$ will model the required variability within the corresponding class. Due to the randomness in the network’s prototypes, the inference of the model will be stochastic, and will be denoted as $f^{\bm{R}}(x)$. In this way, a predictive distribution for the output can be defined:
@@ 3.2 Capturing the predictive uncertainty
The predictive distribution defined in Equation (1) allows us to better model the predictive uncertainty of a Prob-PSENN compared to its deterministic counterpart. First, the total predictive uncertainty can be measured by means of the entropy [36] of the prediction: ${H[p(\bm{y}|x)]=-\sum_{i=1}^{c}p(\bm{y}=y_{i}|x)\cdot\log p(\bm{y}=y_{i}|x)}$· However, for a model with fixed (pointwise) values for the parameters, the source of this uncertainty will be only aleatoric uncertainty, which refers to the possibly unavoidable noise, randomness or ambiguity inherent in the data. Considering a probability distribution for the parameter values instead of point estimates allows us to also capture the epistemic uncertainty [19, 21, 24], which accounts for the model uncertainty about the true parameters modeling the data (e.g., as a consequence of insufficient training data). This enables a safer use of the model in critical tasks, since it allows us to identify when the model is making unreliable predictions. An illustrative comparison between aleatoric and epistemic uncertainty is provided in Figure 2 (left).
Figure 2: Comparison between the predictive and explanatory uncertainty, illustrated for a binary classification problem with a two-dimensional input space. In the left subfigure, dashed lines represent different decision boundaries. In the right subfigure, colored regions represent high density regions for the prototype distribution $p(r_{1})$ and $p(r_{2})$.
Thus, for the predictive distribution defined in Equation (1), $H[p(\bm{y}|x)]$ encapsulates both aleatoric and epistemic uncertainty. To measure epistemic uncertainty only, the mutual information between the output $\bm{y}$ and the parameters $\bm{R}$ can be employed [18, 12, 19, 10], which, in our case, can be formalized as: $I(\bm{y},\bm{R}|x)=H[\,p(\bm{y}|x)]-\mathbb{E}_{p(\bm{R})}\big{[}H[\,p(\bm{y}|x,\bm{R})]\big{]},$ while $\mathbb{E}_{p(\bm{R})}\big{[}H[\,p(\bm{y}|x,\bm{R})]\big{]}$ represents a measure of aleatoric uncertainty. Notice that the values of all these metrics can be normalized to the range $[0,1]$, where $1$ represents the maximum uncertainty, by dividing the resulting value by the entropy of the uniform categorical distribution of $c$ categories. In practice, $H[p(\bm{y}|x)]$ can be approximated by $H[\tilde{p}(\bm{y}|x)]$ (see Equation 2).
@@ 3.3 Capturing the explanation uncertainty
With a PSENN with fixed prototypes, only a point estimate $\delta_{x}$ can be provided for the input at hand, that is, a point estimate on how close the input is to the particular prototypes selected to represent each class, as illustrated in Figure 3-(a). In contrast, the probability distribution over the prototypes considered in Prob-PSENNs induces a probability distribution over each value of $\delta_{x}$, $\,p(\bm{\delta}_{x,i})\!=\!p\bigl{(}\phi(e_{x},\bm{r}_{i})\bigr{)}$, as exemplified in Figure 3-(b) and Figure 3-(c). This allows us to compute not only more representative information for each value of $\delta_{x}$, but also to compute the uncertainty in the explanation, as will be detailed below. Furthermore, this enables a direct and explicit connection between the explanation uncertainty and the output uncertainty, which is a key advantage of Prob-PSENNs.
3.3.1 Types of explanation uncertainty
Similarly to the predictive uncertainty, different types of uncertainty can be attributed to the explanations of Prob-PSENNs. In the context of explanations, we will refer to aleatoric uncertainty as the uncertainty in determining which class-prototypes best represent the input at hand, that is, in determining which prototypes are closest to the encoded input, possibly due to ambiguity in the representation of the input. Therefore, high aleatoric uncertainty will imply highly overlapping distance-distributions $p(\bm{\delta}_{x,i})$. In contrast, epistemic explanatory uncertainty will be high for those inputs that lie far from the training distribution, and, consequently, also far from the distribution over the prototypes, resulting in uninformative and non-representative explanations for that input. Therefore, high epistemic uncertainty will imply large distances with respect to all the prototypes. An illustrative representation of these types of explanatory uncertainty is provided in Figure 2 (right).
3.3.2 Quantifying explanation uncertainty
Given that, in the general case, computing $p(\bm{\delta}_{x,i})$ will be intractable both analytically and numerically, we will rely on sample-based estimators in order to compute meaningful information about those distributions. To begin with, assuming $\delta_{x}^{(1)},\dots,\delta_{x}^{(N)}$ are the distance-vectors obtained for $N$ inferences of the model (i.e., for $N$ random sets of prototypes), the sample mean ${\bar{\delta}_{x,i}=\frac{1}{N}\sum_{n=1}^{N}\delta_{x,i}^{(n)}}$ can inform us about the average similarity between the input and the prototypes of the class $y_{i}$, while the sample variance Var ${\bigl{(}p(\bm{\delta}_{x,i})\bigr{)}\approx\frac{1}{N-1}\sum_{n=1}^{N}\bigl{(}\delta_{x,i}^{(n)}-\bar{\delta}_{x,i}\bigr{)}^{2}}$ can be considered a measure of uncertainty about those similarities.